OpenStudy (anonymous):

Need help finding range of y = 2sin(x-3) with step by steps ? ???

OpenStudy (cwrw238):

the sine function ranges in value from -1 to 1 so 2 sine has range twice that -2 to 2

OpenStudy (cwrw238):

check this link http://www.wolframalpha.com/input/?i=y+%3D+2sin%28x-3%29

OpenStudy (anonymous):

@charlotteakina, does that make sense? It wouldn't matter what you had inside the parenthesis, because that doesn't affect how "big" the sine function curve is. The sine curve normally ranges between -1 and 1, so it you multiply it by 2, you extend its range from -2 to 2. Any coefficient to the sine has the same effect.

OpenStudy (anonymous):

do you have the basic form showing a sine function with a, h, and k? I can't recall how they show it... I can look it up, but if you have it handy, that would help.

OpenStudy (anonymous):

Found it... Is this what you have? y - k = A Sin B(x - h)

OpenStudy (anonymous):

When you get a minute, check this link also... this is what I was using to refresh my memory: http://home.windstream.net/okrebs/page83.html So, the "k" is 0 in your problem... therefore there is no vertical shift.

OpenStudy (anonymous):

y = 2sin(x-3) y - k = A Sin B(x - h) k = 0 <<--- no vertical shift A = 2 <<--- increases range from -1,1 to -2,2 B = 1 <<-- no change in period h = 3 <<--- shifts graph right by 3

OpenStudy (anonymous):

So y = 2sin(x-3) will look like a normal sine curve, except the "ups" and "downs" will be twice as big, and the whole thing will be shifted right by 3 units. If the equation had included a "k" term, like: y - 5 = 2sin(x-3) Then the shape of the curve would be the same as above, but it would be shifted vertically up by 5 units. This would change the range from -2,2 (a "swing" of 4) to 3,7 (still a swing of 4, but shifted up vertically 5 units).

OpenStudy (anonymous):

I'm back. I would simplify it a little. That paragraph is just for your understanding... no need to put all that in, plus I was intentionally trying to talk in "plain terms" vs. "math terms". Why don't you leave out the paragraph and use the other parts...

OpenStudy (anonymous):

The question only asked about range, so it doesn't really matter that it shifts right. But if you have to answer about the A, h, and k parts, then the above explanation does address those aspects. Go eat ;)